driving at 43 mph, one covers 17 miles in how many seconds

Found 2 solutions by Alan3354, richard1234:
Answer by Alan3354(69443)

Here's an example. Do it like this:
driving at 55 mph one covers 19 miles in how many seconds
----------
d = rt
t = d/r
t = 19/55 hours
t = 3600*19/55 seconds
-----------------------
If you have a lot of these, make an Excel sheet.
-----------------
Are you stupid, or lazy, or both?

Answer by richard1234(7193)

You seem to have a lot of these type of problem. I suggest that instead of posting ten of them at once, look at the solution to one of them and be able to generalize and come up with solutions to all of them. Or, you could establish a general formula (e.g. driving at x mph, one covers y miles...) and evaluate for the given numbers.